OpenStudy (anonymous):

find the value of 1+3(5-17)/2*6

OpenStudy (ash2326):

Nicky it's easy if you follow th rules solve the bracket first then division followed by multiplication then addition and last subtraction

OpenStudy (anonymous):

Oh okay, I get it now, I am terrible at math so

OpenStudy (ash2326):

no problem nicky, you'll learn it

OpenStudy (anonymous):

So the answer is 40 correct? Because thats what I got

jhonyy9 (jhonyy9):

1+(15-51)/12 =1+(-36)/12 =1+(-3)=-2

jhonyy9 (jhonyy9):

- so this is the answer than the denominator is (2*6)

jhonyy9 (jhonyy9):

than just 2 is the denominator so than will get 1-18*6=1-108=-107